the probabilities a team wins draws or loses any particular game at 0.6, 0.1, 0.3 respectively. find the probability that the teams wins at least one of its next two games


1
Expert's answer
2022-05-08T14:10:33-0400
"P(W)=p=0.6, P(not\ W)=1-0.6=0.4"

"P(at\ least\ one\ W)=1-P(not\ W\&not\ W)"

"=1-P(not\ W)P(not\ W)=1-0.4(0.4)=0.84"

The probability that the teams wins at least one of its next two games is "0.84."